OpenStudy (anonymous):

How do you solve this ? freaking stressed af A video game has been developed whereby the player must earn a certain number of experience points before progressing to the next level. On level 1, a player must earn 1 experience point. On level 2, a player must earn 6 experience points. On level 3, a player must earn 13 experience points. On level 4, a player must earn 22 experience points. How many experience points must a player earn on level 7?

jimthompson5910 (jim_thompson5910):

from 1 to 6, you have a jump of +5 (since 6-1 = 5) from 6 to 13, you have a jump of +7 (since 13-6 = 7) from 13 to 22, you have a jump of +9 (since 22-13 = 9)

jimthompson5910 (jim_thompson5910):

So the increasing pattern is: 5, 7, 9, ... it keeps going like this 5, 7, 9, 11, 13, 15, ... So Level 5: 22+11 = 33 Level 6: 33+13 = 46 Level 7: 46+15 = 61 So the player must earn 61 points on level 7

OpenStudy (anonymous):

Thank you for your time :D

jimthompson5910 (jim_thompson5910):

yw